OpenStudy (anonymous):

MEDAL FOR RIGHT ANSWER!!! use mental math to calculate 49% of 162

OpenStudy (anonymous):

ok so you know your metric stuff right?

OpenStudy (anonymous):

@mathmath333 @I_Always_Smiling @WOLF__ @rainbowhorse

OpenStudy (anonymous):

79.38 is the answer

OpenStudy (anonymous):

\[162\times.49\] because its 49 over 100 which we write like so \[49/100\] now basically what the / means is 2 two things A its a fraction and B. its division since the slash is used for both

OpenStudy (anonymous):

whats metric stuff?

OpenStudy (anonymous):

@bohotness

OpenStudy (bohotness):

yes

OpenStudy (anonymous):

@kittenlover731 can I have medal the answer is 79.38

OpenStudy (mathmath333):

\(49\%\)is same as \(50\%-1\%\) so \(50\%*162-1\%1*62\\=81-0.62=?\)

OpenStudy (anonymous):

I gave you a medal for you answering half the answer but you still didn't tell me how to use mental math to solve it

OpenStudy (mathmath333):

thats the proper mental math technique i think well \(50\%\) is same as\(\dfrac{1}{2}\) and \(1\%\) is same as \(0.01\) so \(\dfrac{1}{2}*162-0.01*162\)

OpenStudy (mathmath333):

https://www.youtube.com/watch?v=YWHx0RsRNzI

OpenStudy (anonymous):

@ballsyjr

OpenStudy (anonymous):

@wwhitlock

OpenStudy (anonymous):

@quin100

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@littlebird

OpenStudy (anonymous):

@jesstho.-.

OpenStudy (anonymous):

@unklerhaukus

OpenStudy (anonymous):

@marc

OpenStudy (unklerhaukus):

@mathmath333's technique is excellent, (but there was a small mistake) \[49\% \text{ of } 162\\ =49\%\times162\\ =(50\%-1\%)\times162\\ =\left(\frac12-\frac1{100}\right)\times162\\ =\frac{162}2-\frac{162}{100}\\ =81-1.62\\ =80-0.62\\ =79.38\]